If a load of mass is raised up to a height of 15m. What is the work done on the process? (g= 10m/s2)
1
Expert's answer
2020-05-29T10:06:34-0400

The work is done on this process equal to the changing the potential energy of the mass. Suppose, the initial potential energy was "U_0 = 0J". The final potential energy is:


"U = mgh"


where "m" is the mass of the body and "h = 15 m" is the  height  the body was raised to.

Thus, the work will be:


"A = U - U_0 = mgh - 0 = mgh\\
A = m\cdot 10 m/s^2\cdot 15 m = 150m\space J"


where "m" is mass of the body in kg.


Answer. A = 150m J.